# Latent Domain Prompt Learning for Vision-Language Models

### v1

@@ 2.1 Problem setup
Specifically, we consider a Multi-Source Domain Generalization (MSDG) problem with $L$ source domains $S_{l}=\{(x_{i}^{l},y_{i}^{l})\}_{i=1}^{n_{l}}$, each associated with a joint distribution $P_{XY}^{l}$. Note that $P_{XY}^{l}\neq P_{XY}^{l^{\prime}},\forall l,l^{\prime}\in\{1,\cdots,L\}$ and $l\neq l^{\prime}$. The goal of the MSDG problem is to learn a predictive function $f:\mathcal{X}\to\mathcal{Y}$ using source domain data such that it minimizes the prediction error on an unseen target domain $S_{target},P_{XY}^{target}\neq P_{XY}^{l},\forall l\in\{1,\cdots,L\}$:
@@ 2.2 Soft prompt design
In our settings, each prompt is divided into three parts: domain-agnostic, domain-specific, and class label, as follows:
@@ 2.3 Latent domain model
where $h(\cdot)$ represents the auxiliary classifier, and $f(\cdot)$ denotes the image encoder. The resulting domain features are then clustered using k-means, which assigns each sample a latent domain label. The cluster centroids are stored and later used during inference to compute similarities between input images and latent domains.
@@ 2.4 Prompt fusion mechanism
We represent the text features from different domains as a set of basis vectors and obtain the fused feature for class $i$ via a weighted linear combination:

### v2

@@ 2.1 Problem setup
Specifically, we consider a Multi-Source Domain Generalization (MSDG) problem with $L$ source domains $S_{l}=\{(x_{i}^{l},y_{i}^{l})\}_{i=1}^{n_{l}}$, each associated with a joint distribution $P_{XY}^{l}$. Note that $P_{XY}^{l}\neq P_{XY}^{l^{\prime}},\forall l,l^{\prime}\in\{1,\cdots,L\}$ and $l\neq l^{\prime}$. The goal of the MSDG problem is to learn a predictive function $f:\mathcal{X}\to\mathcal{Y}$ using source domain data such that it minimizes the prediction error on an unseen target domain $S_{target},P_{XY}^{target}\neq P_{XY}^{l},\forall l\in\{1,\cdots,L\}$ [16]:
@@ 2.2 Soft prompt design
In our settings, each prompt is divided into three parts: domain-agnostic, domain-specific, and class label:
@@ 2.3 Latent domain model
where $h(\cdot)$ represents the auxiliary classifier, and $f(\cdot)$ denotes the image encoder. The resulting domain features are then clustered using k-means, which assigns each sample a latent domain label. We set the number of clusters to $k=N_{s}$, a hyperparameter tuned via the validation set, strictly independent of ground-truth domain labels. The cluster centroids are stored and later used during inference to compute similarities between input images and latent domains.
@@ 2.4 Prompt fusion mechanism
This mechanism generates an adaptive text feature that aligns with the target image despite domain shift. We represent the text features from different domains as a set of basis vectors and obtain the fused feature for class $i$ via a weighted linear combination: